rewrite the equation 2x-3y=-6 as a function of x
Answer by tutorcecilia(2152)

2x-3y=-6 [solve for the y-value]
2x-2x-3y=-2x-6
-3y=-2x-6
-3y/-3=-2x/-3-6/-3
y=2/3x+2
f(x)=2/3x+2 [re-write as a function of x]